### v4

# Parallel Split Learning with Global Sampling

Abstract
Distributed deep learning in resource-constrained environments faces scalability and generalization challenges due to large effective batch sizes and non-identically distributed client data. We introduce a server-driven sampling strategy that maintains a fixed global batch size by dynamically adjusting client-side batch sizes. This decouples the effective batch size from the number of participating devices and ensures that global batches better reflect the overall data distribution. Using standard concentration bounds, we establish tighter deviation guarantees compared to existing approaches. Empirical results on a benchmark dataset confirm that the proposed method improves model accuracy, training efficiency, and convergence stability, offering a scalable solution for learning at the network edge.
Index Terms:
Parallel split learning, distributed deep learning, non-iid data, IoT.
I Introduction
Distributed deep learning (DDL) has emerged as a powerful paradigm in the era of the Internet of Things (IoT), addressing the challenges posed by the proliferation of resource-constrained devices and the exponential growth of data at the network edge [1, 2, 3]. Split learning (SL) [4, 5] is a DDL paradigm designed for environments with constrained computational and communication resources, such as IoT settings. In SL, neural network training is partitioned between client devices and a central server. This setup enhances data privacy, minimizes client-side computation, and facilitates faster convergence [4, 6, 7]. However, the sequential execution across devices introduces considerable training and communication latency [6].
To address training latency in SL, parallel split learning (PSL) [8, 9, 10, 11, 12] has been proposed and introduces parallelism without the need for multiple server model instances like in split federated learning (v1) [13] or client model aggregation like in FedAvg [14]. However, PSL encounters the large effective batch size problem [15, 11], where the effective batch size scales with the number of clients. Additionally, devices usually contribute from local datasets that are not independent and identically distributed (IID) and vary in sizes, causing batch distributions to deviate from the overall data distribution, which can lead to poor generalization. This issue of non-IID data is a general problem in DDL, not specific to PSL.
Figure 1: Schematic illustration of the proposed GPSL method with variable local batch sizes summing to a fixed global batch size, mitigating the large effective batch size and non-IID data problems in PSL.
We propose parallel split learning with global sampling (GPSL) building upon the original PSL framework introduced in [8] to address the large effective batch size problem and the issue of non-IID data. GPSL allows for variable local batch sizes that sum up to a fixed global batch size. This decouples the number of clients from the effective batch size. We show that this decoupling is crucial for controlling the deviation probability of global batches from the overall data distribution, leading to improved model performance and reduced training latency in non-IID settings. In particular, we show that the choice of local batch sizes is crucial for controlling the deviation probability. An overview of GPSL is shown in Figure 1. The main contributions of this paper are listed below:
  • We introduce the GPSL method to address the large effective batch size problem and the non-IID data problem in PSL.
  • We analyze the global batch deviation and establish tight deviation bounds for GPSL.
  • Our simulation results show that GPSL reduces the deviation probability of global batches significantly, leading to improved accuracy and reducing training latency.
The remainder of this paper is organized as follows. Sections II and III present related works and preliminaries of our work. In Section IV, we introduce our novel GPSL method and analyze the deviation probability of global batches. Section V presents the simulation results and addresses known limitations. Section VI concludes the paper with final remarks. Code and results available on GitHub [16].
II Related Work
Jeon and Kim [8] introduce PSL and use gradient averaging over all clients to address the client decoupling problem, where the information flow between clients is decoupled during BP. They fixed the local batch sizes proportional to the dataset size of each client. Lyu et al. [17] propose averaging the clients’ local loss function outputs instead of their local gradients. This approach aligns with that of [8] due to the linearity of the gradient operator, but enhances privacy as clients can compute the local loss functions without sharing labels. Both show stable convergence and improved accuracy in both IID and non-IID settings. However, the proportion of each client’s data in the total was chosen to yield an integer when scaled by the global batch size to determine local batch sizes. In practice, rounding is typically necessary when determining each client’s local batch size, which can lead to imbalances in the global batch composition. Furthermore, the large effective batch size problem is not addressed in these works.
To mitigate the large effective batch size problem, Pal et al. [15] propose scaling the learning rate on the server side while keeping the local batch size fixed and identical for all clients. Following the line of work from [12], Oh et al. [11] propose using smashed data augmentation, local loss function design and FedAvg for the client models to overcome the large effective batch size problem and client model decoupling. Pal et al. [15] and Oh et al. [11] assume that the effective batch size scales with the number of clients. However, it has been shown that increasing the batch size leads to diminishing returns for gradient estimation [18] and that smaller batch sizes [19] can be more beneficial for model generalization, if communication constraints allow. Furthermore, the empirical results are reported only for IID settings.
Cai and Wei [20] propose an incentive mechanism to handle non-IID data in PSL, which encourages clients with larger and more evenly distributed datasets to participate in training. The bargaining game approach for personalized SL [21] balances energy consumption and privacy in non-IID settings. However, these methods may face scalability issues and may not fully address the heterogeneity of data distributions across diverse clients.
III Preliminaries
We consider the PSL setting from [8]. The client index set is $\mathcal{K}=\{1,2,\dots,K\}$. A server and $|\mathcal{K}|=K$ clients train a neural network model with parameter vector $\boldsymbol{w}$ collaboratively using the PSL framework. We focus on supervised classification tasks, where the dataset $\mathcal{D}_{k}$ of each client $k\in\mathcal{K}$ consists of input-label pairs $\mathcal{D}_{k}=\{(\boldsymbol{x}_{k,i},y_{k,i})\}_{i=1}^{D_{k}}$, where $D_{k}\in\mathbb{N}$ is the dataset size, $\boldsymbol{x}_{k,i}$ is the input data, and $y_{k,i}\in\mathcal{M}$ is the ground truth class label of $\boldsymbol{x}_{k,i}$. Here, =$\mathcal{M}=\{1,2,\dots,M\}$ represents all classes and $M\in\mathbb{N}$ is the number of classes. The overall dataset $\mathcal{D}_{0}=\bigcup_{k\in\mathcal{K}}\mathcal{D}_{k}$ is the union of all clients’ datasets, with a total size $D_{0}=\sum_{k\in\mathcal{K}}D_{k}$.
The class distribution of a client $k$ with class label vector $\boldsymbol{y}_{k}\in\mathcal{M}^{D_{k}}$ is denoted as $\boldsymbol{\beta}_{k}\in\mathbb{R}^{M}$ with
${\boldsymbol{\beta}_{k}=\frac{1}{D_{k}}(c(\boldsymbol{y}_{k},1),c(\boldsymbol{y}_{k},2),\dots,c(\boldsymbol{y}_{k},M))^{\top}},$ (1)
where $c(\boldsymbol{y},m)$ counts the number of times class $m$ appears in $\boldsymbol{y}$. The overall class distribution of $\mathcal{D}_{0}$ with class label vector $\boldsymbol{y}_{0}\in\mathcal{M}^{D_{0}}$ is denoted as $\boldsymbol{\beta}_{0}\in\mathbb{R}^{M}$ according to the same definition.
Every epoch consists of $T\in\mathbb{N}$ optimization steps. The set of optimization steps is denoted as $\mathcal{T}=\{1,2,\ldots,T\}$. Each client $k$ has a local batch size $B_{k}^{(t)}\in\mathbb{N}$. The global batch size $B^{(t)}=\sum_{k\in\mathcal{K}}B_{k}^{(t)}$ is the total number of intermediate representations that arrive at the server in each optimization step $t\in\mathcal{T}$. Accordingly, we differentiate between local batches $\mathcal{B}^{(t)}_{k}\subseteq\mathcal{D}_{k}$ and the global batch $\mathcal{B}^{(t)}=\bigcup_{k\in\mathcal{K}}\mathcal{B}^{(t)}_{k}$.
PSL splits the model into client-side parameters $\boldsymbol{w}_{c}$ and server-side parameters $\boldsymbol{w}_{s}$. Before training, the server sends $\boldsymbol{w}_{c}$ to the clients, each storing a copy $\boldsymbol{w}_{c,k}$ with identical initial values. Accordingly, the model function $f(\boldsymbol{x};\boldsymbol{w})$ is split into two functions $f_{c}$ and $f_{s}$ as $f(\boldsymbol{x};\boldsymbol{w})=f_{s}(f_{c}(\boldsymbol{x};\boldsymbol{w}_{c});\boldsymbol{w}_{s})$. The goal is to find the optimal model parameters ${\boldsymbol{w}^{*}=(\boldsymbol{w}_{s}^{*},\boldsymbol{w}_{c}^{*})^{\top}}$ that minimize the global loss function. Each optimization step $t\in\mathcal{T}$ involves the following substeps:
  1. 1.
    Parallel Client-Side Forward Propagation (FP): Each client $k$ samples a local batch $\mathcal{B}^{(t)}_{k}\subseteq\mathcal{D}_{k}$ with a local batch size of $B_{k}^{(t)}$ and sends $f_{c}(\mathcal{B}^{(t)}_{k};\boldsymbol{w}_{c,k})$ with the corresponding labels to the server.
  2. 2.
    Server-Side FP: The server concatenates the received activations to obtain the global batch ${\mathcal{B}^{(t)}=\bigcup_{k\in\mathcal{K}}f(\mathcal{B}^{(t)}_{k};\boldsymbol{w}_{c,k})}$ and computes the final predictions $f_{s}(\mathcal{B}^{(t)};\boldsymbol{w}_{s})$ to compute the average loss.
  3. 3.
    Server-Side Backward Propagation (BP): The server computes the server-gradients through backpropagation, updates its model parameters according to the optimization algorithm, and sends the gradient of the first server-side layer (cut layer) to all clients.
  4. 4.
    Parallel Client-Side BP: Each client $k$ continues backpropagation using the cut layer gradient to compute the client-gradients $\boldsymbol{g}^{(t)}_{\boldsymbol{w}_{c,k}}$ and sends it back to the server.
  5. 5.
    Server-Side Gradient Averaging: The server computes the average client gradients $\boldsymbol{\overline{g}}^{(t)}_{\boldsymbol{w}_{c}}=\sum_{k\in\mathcal{K}}\frac{D_{k}}{D_{0}}\boldsymbol{g}^{(t)}_{\boldsymbol{w}_{c,k}}$ and sends it to all clients.
  6. 6.
    Parallel Client-Side Model Update: Each client uses $\boldsymbol{\overline{g}}^{(t)}_{\boldsymbol{w}_{c}}$ to update its model parameters.
An epoch is concluded after $T$ optimization steps. $T$ is determined by the number of non-empty global batches that can be formed from the clients’ datasets until depletion and thus depends on the sampling procedure. The model is trained for a fixed number of epochs $E\in\mathbb{N}$.
IV System Model
This section introduces the proposed GPSL method and analyzes how it improves the alignment between global batches and the overall data distribution compared to random uniform sampling in a centralized setting.
IV-A Parallel Split Learning with Global Sampling
In GPSL, we replace the notion of fixed per‐client batch sizes with a single global batch size $B$. The server first gathers each client’s dataset size $D_{k}$. This information can trivially be obtained from the server through counting the number of samples that arrive from each client and is not sensitive. Before training the server must choose how many samples $B_{k}^{(t)}$ each client $k$ will contribute to each training step $t\in\mathcal{T}$, subject to
$\sum_{k\in\mathcal{K}}B_{k}^{(t)}\;=\;B.$
Rather than uniformly fixing these $B_{k}^{(t)}$, we draw the local batch sizes using a global sampling procedure over the pooled data. We first set each client’s remaining dataset size $D_{k}^{(0)}=D_{k}$. Then, for each $t\in\mathcal{T}$:
  1. 1.
    Dynamic Sampling. For $i=1,\dots,B$:
    1. (a)
      Compute client weights
      $\pi_{k}^{(i)}=D_{k}^{(i)}/\sum_{j\in\mathcal{K}}D_{j}^{(i)}.$ (2)
    2. (b)
      Sample a client index $z_{i}\sim\mathrm{Categorical}(\boldsymbol{\pi}^{(i)})$.
    3. (c)
      Decrement its remaining dataset size
      $D_{z_{i}}^{(i+1)}\leftarrow D_{z_{i}}^{(i)}-1.$ (3)
  2. 2.
    Batch Composition. After $B$ draws, the number of samples from client $k$ are obtained by
    $B_{k}^{(t)}\;=\;|\{\,i:z_{i}=k\}|.$ (4)
    Remaining dataset sizes are reset to $D_{k}^{(0)}=D_{k}^{(i)}$
The server forwards the $\{B_{k}^{(t)}\}^{t\in\mathcal{T}}$ to each client before training starts. During training clients uniformly sample random mini-batches using $B_{k}^{(t)}$ as local batch size. Because each draw in step 1) selects one of the remaining points uniformly without replacement, the resulting batch during training is exactly a uniform draw at random of $B$ distinct samples from the union $\mathcal{D}_{0}$. The rest of the PSL training loop remains unchanged, allowing seamless integration of the sampling procedure into existing PSL frameworks.
IV-B Deviation Analysis
We assess how well a batch $\mathcal{B}\subset\mathcal{D}_{0}$ of size $B$ reflects the true label distribution $\boldsymbol{\beta}_{0}$ using the $\ell_{1}$ deviation:
$\Delta(\mathcal{B},\boldsymbol{\beta}_{0})=\sum_{m\in\mathcal{M}}\bigl{|}\hat{\beta}_{m}-\beta_{0,m}\bigr{|},\quad\hat{\beta}_{m}=\frac{1}{B}\,c(\mathcal{B},m).$ (5)
[…]
$B\leq D_{0},\quad 0<\varepsilon\leq 1-\frac{1}{D_{0}},$ (6)
so that all finite‐population corrections remain positive and $\varepsilon B\leq B-1$. When $\mathcal{B}$ is drawn by sampling $B$ points without replacement from $\mathcal{D}_{0}$, the count $Y_{m}=c(\mathcal{B},m)$ follows a hypergeometric distribution
$Y_{m}\sim\mathrm{Hypergeometric}(D_{0},\,c(D_{0},m),\,B).$
with $\mathbb{E}[Y_{m}]=B\,\beta_{0,m}$. By Serfling’s inequality with finite‐population correction [22],
$\Pr\bigl{(}|Y_{m}-B\beta_{0,m}|\geq\varepsilon B\bigr{)}$ $=\Pr\!\Bigl{(}\bigl{|}\tfrac{Y_{m}}{B}-\beta_{0,m}\bigr{|}\geq\varepsilon\Bigr{)}$ (7)
[…]
Although the $\{Y_{m}\}^{m\in\mathcal{M}}$ are dependent as they are required to sum up to $B$, a union bound yields
$\Pr\!\Bigl{(}\Delta(\mathcal{B},\boldsymbol{\beta}_{0})\geq M\varepsilon\Bigr{)}$ $\leq\sum_{m\in\mathcal{M}}2\exp\!\Bigl{(}-\tfrac{2\,\varepsilon^{2}\,B}{1-\tfrac{B-1}{D_{0}}}\Bigr{)}$ (9)
$=2M\exp\!\Bigl{(}-\tfrac{2\,\varepsilon^{2}\,B}{1-\tfrac{B-1}{D_{0}}}\Bigr{)}.$ (10)
By construction, GPSL’s depletion‐aware sampling is exactly equivalent to drawing $B$ points uniformly at random without replacement from the pooled dataset $\mathcal{D}_{0}$. Consequently, the class counts $Y_{m}$ in GPSL follow the same hypergeometric law as in CL. Applying Serfling’s inequality with finite‐population correction and then a union bound over $\mathcal{M}$ gives
$\Pr\!\Bigl{(}\Delta(\mathcal{B},\boldsymbol{\beta}_{0})\geq M\varepsilon\Bigr{)}$ $=\sum_{m\in\mathcal{M}}\Pr\!\Bigl{(}\bigl{|}\tfrac{Y_{m}}{B}-\beta_{0,m}\bigr{|}\geq\varepsilon\Bigr{)}$ (11)
$\leq 2M\exp\!\Bigl{(}-\frac{2\,\varepsilon^{2}\,B}{1-\tfrac{B-1}{D_{0}}}\Bigr{)}.$ (12)
Thus GPSL attains the exact same exponential‐tail deviation guarantee (up to the standard finite‐population factor) as centralized uniform sampling.
IV-C Comparison with Fixed Local Batches
Under a scheme with a fixed local batch size as in [8], each client $k$ contributes $\tilde{B}_{k}=\lceil B\cdot\tfrac{D_{k}}{D_{0}}\rceil$ samples without replacement in each step. Defining $\tilde{Y}^{\prime}_{m}=\sum_{k\in\mathcal{K}}\tilde{Y}_{k,m}$ with
$\tilde{Y}_{k,m}\sim\mathrm{Hypergeometric}(D_{k},c(D_{k},m),\tilde{B}_{k})$
and $\tilde{p}_{m}=\sum_{k\in\mathcal{K}}\tfrac{\tilde{B}_{k}}{B}\,\beta_{k,m}$ we decompose
$\bigl{|}\tfrac{\tilde{Y}^{\prime}_{m}}{B}-\beta_{0,m}\bigr{|}\;\leq\;\underbrace{\bigl{|}\tfrac{\tilde{Y}^{\prime}_{m}}{B}-\tilde{p}_{m}\bigr{|}}_{\text{sampling noise}}\;+\;\underbrace{\bigl{|}\tilde{p}_{m}-\beta_{0,m}\bigr{|}}_{\text{rounding bias}}.$ (13)
The sampling noise term still decays exponentially, but the rounding bias satisfies
$\bigl{|}\tilde{p}_{m}-\beta_{0,m}\bigr{|}$ $=\Bigl{|}\sum_{k\in\mathcal{K}}\Bigl{(}\tfrac{\tilde{B}_{k}}{B}-\tfrac{D_{k}}{D_{0}}\Bigr{)}\beta_{k,m}\Bigr{|}$ (14)
$\leq\;\sum_{k\in\mathcal{K}}\Bigl{|}\tfrac{\tilde{B}_{k}}{B}-\tfrac{D_{k}}{D_{0}}\Bigr{|}=O\!\bigl{(}\tfrac{K}{B}\bigr{)}.$ (15)
When $K/B$ is large, the rounding bias can dominate and renders any exponential‐tail guarantee ineffective. By contrast, GPSL’s server‐driven batch allocation incurs zero per‐class rounding bias while retaining the same exponential decay. Fixed local batches may still perform adequately when $K\ll B$ and rounding effects are negligible. However, such settings are rare in edge learning environments, where large $K$ and small batch sizes dominate due to memory and communication constraints.
Moreover, the rounding exacerbates the issue of client data depletion, where client datasets are depleted disproportionately and faster. This increases the total number of training steps $T$ and prolongs training. By contrast, GPSL prevents unnecessary inflation of $T$ by keeping the global batch size $B$ constant, regardless of the number of clients.
V Simulation Results
[…]
V-A Sampling Methods
We compare GPSL against two sampling methods commonly used in PSL:
  • Fixed local sampling (FLS) [15]: Each client $k$ has the same local batch size $B_{k}=\lceil B/K\rceil$.
  • Fixed proportional sampling (FPLS) [8]: Each client $k$ has a proportional $B_{k}\propto D_{k}$ with $B_{k}=\lceil B\cdot\frac{D_{k}}{D_{0}}\rceil$.
We also include a centralized learning (CL) baseline using random uniform sampling, where all data are pooled at the server and trained with a single model. This serves as an upper bound for performance.
In Table III we report the results of the sampling methods in the IID setting, where all clients have the same distribution. All sampling methods perform similarly, achieving near central learning performance.
V-B Non-IID Data
In the non-IID scenario, we employ a modified version of the extended Dirichlet allocation method proposed in [7], which ensures that each client receives data from exactly $C$ classes. First, $C$ distinct classes are assigned to each client using a multinomial-based allocation that promotes balanced class coverage across the client population. Then, for each class, a Dirichlet distribution with concentration parameter $\alpha$ is sampled to determine how its samples are distributed among the clients assigned to it. Lower values of $\alpha$ lead to more imbalanced allocations, where a class is concentrated heavily on one or a few clients. Higher values of $\alpha$ result in more uniform class sample distributions across clients. By tuning $C$ and $\alpha$, we can control the degree and structure of statistical heterogeneity across clients. We consider a mild non-IID scenario with $C=5$ and $\alpha=3.0$ and a severe non-IID scenario with $C=2$ and $\alpha=3.0$. In Figure 2 we show an example of the data distribution for $K=16$ clients in the severe non-IID scenario.
Figure 2: Example of a severe non-IID setting for $K=16$ clients, showing the distribution of data across classes.
TABLE I: IID test accuracy on CIFAR10
Method $\mathbf{K=16}$ $\mathbf{K=32}$ $\mathbf{K=64}$ $\mathbf{K=128}$
FLS 84.21 $\pm$ 0.50 84.31 $\pm$ 0.54 84.15 $\pm$ 0.79 84.10 $\pm$ 0.36
FPLS 84.42 $\pm$ 0.58 84.31 $\pm$ 0.54 84.15 $\pm$ 0.79 84.10 $\pm$ 0.36
GPSL 84.74 $\pm$ 0.47 84.40 $\pm$ 0.56 84.29 $\pm$ 0.53 84.66 $\pm$ 0.45
TABLE II: Severe Non-IID test accuracy on CIFAR10
Method $\mathbf{K=16}$ $\mathbf{K=32}$ $\mathbf{K=64}$ $\mathbf{K=128}$
FLS 61.55 $\pm$ 7.18 57.17 $\pm$ 4.33 66.02 $\pm$ 6.00 67.73 $\pm$ 9.29
FPLS 59.09 $\pm$ 8.25 64.88 $\pm$ 9.34 68.40 $\pm$ 8.08 67.11 $\pm$ 8.39
GPSL 84.71 $\pm$ 0.29 84.28 $\pm$ 0.20 84.30 $\pm$ 0.32 84.40 $\pm$ 0.51
TABLE III: Mild Non-IID test accuracy on CIFAR10
Method $\mathbf{K=16}$ $\mathbf{K=32}$ $\mathbf{K=64}$ $\mathbf{K=128}$
FLS 66.12 $\pm$ 6.92 70.97 $\pm$ 5.32 70.45 $\pm$ 6.87 69.33 $\pm$ 10.62
FPLS 59.18 $\pm$ 14.48 64.61 $\pm$ 10.05 71.43 $\pm$ 4.92 71.25 $\pm$ 6.08
GPSL 84.50 $\pm$ 0.56 84.34 $\pm$ 0.34 84.52 $\pm$ 0.24 84.43 $\pm$ 0.26
Top-1 test accuracy on CIFAR-10 under IID and non-IID settings for different numbers of clients $K$ and global batch size $B=128$. Each result is averaged over 5 random seeds and reported with standard deviation. CL achieved 84.06±0.5384.06$\pm$ 0.5384.06 ± 0.53. Bold values indicate the best result per column.
Figure 3: Test accuracy curves for each sampling method over 5 runs with standard deviation shaded. Severe non-IID setting with $K=64$ clients and a global batch size of $B=128$.
Table III shows the performance of each sampling method under severe non-IID conditions. Table III shows the performance in a mild non-IID scenario. In both cases, GPSL achieves a test accuracy on par with centralized learning, while FPLS and FLS degrade significantly. The results are consistent across both mild and severe non-IID scenarios for varying client counts $K$, indicating scalability and robustness to the degree of data heterogeneity. Figure 3 shows the test accuracy curves for each sampling method under severe non-IID conditions. GPSL maintains a stable convergence curve, while FPLS and FLS exhibit significant fluctuations.
TABLE IV: Effect of Global Batch Size on Severe Non-IID Test Accuracy
Method $\mathbf{B=64}$ $\mathbf{B=128}$ $\mathbf{B=256}$
FLS 69.57 $\pm$ 2.93 67.73 $\pm$ 9.29 64.39 $\pm$ 4.80
FPLS 69.25 $\pm$ 4.96 67.11 $\pm$ 8.39 64.25 $\pm$ 11.19
GPSL 86.06 $\pm$ 0.28 84.40 $\pm$ 0.51 82.22 $\pm$ 0.45
CL 86.29 $\pm$ 0.24 84.06 $\pm$ 0.53 81.47 $\pm$ 0.80
Top-1 test accuracy on CIFAR-10 under severe non-IID settings for different global batch sizes $B$. Each result is averaged over 5 random seeds and reported with standard deviation. $K=128$ clients. Bold values indicate the best result per column.
V-C Global Batch Size
Table IV shows performance results for the different sampling methods with varying batch sizes $B$. GPSL consistently outperforms FPLS and FLS across all batch sizes, demonstrating its robustness to batch size variations. In particular, GPSL enables flexibility in choosing the global batch size whereas FPLS and FLS are sensitive to the choice of batch size and may result in much larger global batch sizes than intended due to rounding and the client data depletion problem (Section IV-C).
V-D Batch Deviation
Figure 4: Batch deviation curves for each sampling method over 5 runs with exponential moving average smoothing. Severe non-IID setting with $K=64$ clients and a global batch size of $B=128$.
Figure 4 shows the empirical evaluation of the batch deviation as defined in Section IV-B for each sampling method under severe non-IID conditions. GPSL maintains a low and stable batch deviation almost matching CL, while FPLS and FLS exhibit significant fluctuations, particularly in the standard deviation. This aligns with our theoretical analysis and performance results, indicating that the batch deviation is the main reason for the performance degradation of FPLS and FLS.
Note that FLS and FPLS have more training steps than GPSL. This is due to the client data depletion problem, which leads to a larger number of batches per epoch with uneven batch sizes.
V-E Runtime
Figure 5: Average total training time for each sampling method over 5 runs, measured in minutes. Severe non-IID setting with $K=64$ clients.
Figure 5 shows average training time for different global batch sizes $B$. GPSL consistently achieves a lower training time compared to FPLS and FLS. In particular when the global batch size is smaller. The computational overhead introduced by GPSL is thus negligible compared to the total training time saved. This result highlights that local sampling strategies can lead to a significant increase in the total training time due to the client data depletion problem. In particular, when the global batch size is smaller than the total number of clients.
V-F Limitations
While GPSL improves model performance and training efficiency, it inherits the PSL framework’s reliance on gradient averaging, which increases training latency and communication overhead. This limits its applicability to real-time and unstable IoT environments. Future work should explore integrating GPSL with client clustering [27] or communication-efficient techniques, such as adaptive compression [10], to improve scalability and reduce overhead.
VI Conclusion
We introduced parallel split learning with global sampling, a method designed to address the large effective batch size and non-independent and identically distributed data problems in distributed deep learning. By dynamically adjusting local batch sizes while maintaining a fixed global batch size, the method decouples the global batch size from the number of participating clients. Using standard concentration bounds, we showed that global sampling improves the alignment between global batches and the overall data distribution. Our simulation results confirm that the proposed method enhances convergence stability, boosts classification accuracy, and reduces training latency in non-independent and identically distributed scenarios. These findings suggest that server-driven global sampling offers a simple yet effective improvement for scalable distributed learning at the network edge.